### v2

# Refusing Safe Prompts for Multi-modal Large Language Models

Abstract
Multimodal large language models (MLLMs) have become the cornerstone of today’s generative AI ecosystem, sparking intense competition among tech giants and startups. In particular, an MLLM generates a text response given a prompt consisting of an image and a question. While state-of-the-art MLLMs use safety filters and alignment techniques to refuse unsafe prompts, in this work, we introduce MLLM-Refusal, the first method that induces refusals for safe prompts. In particular, our MLLM-Refusal optimizes a nearly-imperceptible refusal perturbation and adds it to an image, causing target MLLMs to likely refuse a safe prompt containing the perturbed image and a safe question. Specifically, we formulate MLLM-Refusal as a constrained optimization problem and propose an algorithm to solve it. Our method offers competitive advantages for MLLM model providers by potentially disrupting user experiences of competing MLLMs, since competing MLLM’s users will receive unexpected refusals when they unwittingly use these perturbed images in their prompts. We evaluate MLLM-Refusal on four MLLMs across four datasets, demonstrating its effectiveness in causing competing MLLMs to refuse safe prompts while not affecting non-competing MLLMs. Furthermore, we explore three potential countermeasures—adding Gaussian noise, DiffPure, and adversarial training. Our results show that though they can mitigate MLLM-Refusal’s effectiveness, they also sacrifice the accuracy and/or efficiency of the competing MLLM. The code is available at https://github.com/Sadcardation/MLLM-Refusal.
1 Introduction
Multimodal large language models (MLLMs) [41, 47, 31, 60, 15, 6] have emerged as a groundbreaking foundation for various generative AI applications such as visual question answering [31], image captioning [26], and embodied AI [17]. MLLMs are typically trained and deployed as web chatbots or cloud API services by resourceful entities, including major technology companies and startups. These companies are fiercely competing in the development of MLLMs, exemplified by Google’s Gemini Pro [47] and OpenAI’s GPT-4 [1]. An MLLM generally comprises three key components: a vision encoder, a vision-language projector, and a large language model (LLM). The vision encoder generates an embedding vector for an image, which the vision-language projector converts into tokens in the LLM’s input token space. These tokens are concatenated with the question’s tokens and fed into the LLM, producing the final text response.
Despite rapid progress in MLLMs’ capabilities, their safety has garnered increasing attention. A recent U.S. Executive Order [54] underscores the commitments from AI companies to ensure the safety and security of advanced AI systems. A safeguarded MLLM typically prevents generating harmful outputs by refusing unsafe prompts [24, 39, 43, 46]. Specifically, if either the prompt’s image or question is unsafe, i.e., containing harmful content, the prompt is considered unsafe. For example, the first two rows in Figure 1 illustrate two types of unsafe prompts and the corresponding refusals from an MLLM. MLLMs can achieve this through safety filters [24, 57] and alignment techniques [43, 46, 13].
Figure 1: Three types of refuals in MLLMs.
Recent studies [44, 36, 49] have shown that an attacker can bypass an MLLM’s refusal capability against unsafe prompts even if the MLLM is safeguarded. These attacks, known as jailbreak attacks, involve adding perturbations to the image [44, 36, 49] or the question [36] to cause an aligned MLLM to generate harmful responses for unsafe prompts. For example, Qi et al. [44] demonstrated that an attacker can optimize an image perturbation, when added to an image, can cause a victim MLLM to generate unsafe responses for unsafe prompts. While existing studies mainly focus on enhancing or bypassing MLLMs’ refusal capabilities against unsafe prompts, an equally critical aspect has remained overlooked: an MLLM’s refusal capability against safe prompts.
Our Work:  In this work, we introduce a novel perspective on refusal: the refusal against safe prompts by MLLMs. A prompt is considered safe only if both the image and the question are devoid of harmful content. The third row in Figure 1 illustrates this concept. We explore scenarios where an image, seemingly benign, is subtly perturbed with a nearly-imperceptible perturbation named refusal perturbation, while the question remains safe. The image and question form a safe prompt that is likely to provoke an unwarranted refusal response from the target MLLM.
We consider three key roles in our setting: an MLLM model provider, its competitors, and normal users. The competitors are also MLLM model providers who develop competing MLLMs. Normal users query MLLMs with safe prompts. We consider the MLLM model provider is the attacker who aims to gain competitive advantages by utilizing effective refusal perturbations to cause competing MLLMs to refuse safe prompts. The model provider perturbs images and publishes them on the Internet, e.g., social media. Normal users of competing MLLMs would experience unexpected refusals when they unwittingly use these perturbed images in their prompts, leading to frustration and a decline in user satisfaction. Such new angle of refusal opens avenues for competitive differentiation in the rapidly evolving field of MLLMs.
MLLM-Refusal. We propose MLLM-Refusal, which optimizes a refusal perturbation to make competing MLLMs refuse safe prompts. Specifically, we consider three primary goals when crafting a refusal perturbation: effectiveness, locality, and stealthiness. Roughly speaking, the effectiveness goal ensures that refusal perturbations cause competing MLLMs to refuse safe prompts containing the refusal perturbation. The locality goal ensures that refusal perturbations are effective only against competing MLLMs, while being ineffective against the model provider’s own MLLM. The stealthiness goal ensures that refusal perturbations are nearly-imperceptible.
Figure 2: Images without and with refusal perturbations added by our MLLM-Refusal under different $\ell_{\infty}$-norm constraint $\epsilon$.
To achieve the three goals, our MLLM-Refusal uses a set of shadow questions to mimic the actual questions of normal users. Given an image, a set of shadow questions, and some competing MLLMs, MLLM-Refusal optimizes a refusal perturbation so that competing MLLMs are likely to refuse safe prompts containing any shadow question and the image embedded with the refusal perturbation. To ensure stealthiness, MLLM-Refusal applies an $\ell_{\infty}$-norm constraint on the refusal perturbation during optimization. Formally, we formulate this as a constrained optimization problem and solve it via a gradient-based method. Figure 2 shows several examples of images with refusal perturbations. Our intuition is that competing MLLMs are likely to refuse users’ safe prompts even when user questions differ from shadow questions, similar to the transferability of adversarial examples [52]. Because different MLLMs have unique vision-language projectors and the refusal perturbation is not optimized for the model provider’s own MLLM, it does not affect the model provider’s own MLLM, thus achieving the locality goal.
Evaluations. Our evaluation of MLLM-Refusal involves using image-question pairs from VQAv2 [3], GQA [22], TextVQA [50], and an extended CelebA [34] dataset. We evaluate MLLM-Refusal’s performance on four open-source MLLMs: LLaVA-1.5 [31], MiniGPT-4 [60], Qwen-VL-Chat [6], and InstructBLIP [15]. We use the refusal rate as our evaluation metric, defined as the proportion of prompts that result in refusals when perturbed images are presented to the MLLMs. MLLM-Refusal applies $\ell_{\infty}$-norm constraint $\epsilon=8/255$ and thus achieves the stealthiness goal since such constraint is considered visually stealthy in previous work [44, 35, 7]. MLLM-Refusal achieves the effectiveness goal across all MLLMs and datasets. For example, MLLM-Refusal achieves over 0.88 refusal rates with LLaVA-1.5 on VQAv2 with three types of shadow questions. Moreover, our MLLM-Refusal maintains near-zero refusal rates on non-competing MLLMs, thereby achieving the locality goal.
Countermeasures. To counter the refusal perturbations created by MLLM-Refusal, we evaluate three countermeasures: adding Gaussian noise, DiffPure [40], and adversarial training. Gaussian noise reduces MLLM-Refusal’s effectiveness but significantly lowers MLLM accuracy, from 0.92 to around 0.80 with standard deviation $\sigma$=0.02. DiffPure uses a diffusion model to purify images, reducing MLLM-Refusal’s refusal rates but also dropping accuracy from 0.92 to 0.78 and increasing inference time by up to 13.07%. Adversarial training, which involves fine-tuning an MLLM on images with refusal perturbations, reduces refusal rates to around 60% but also significantly decreases MLLM accuracy and requires substantial computational resources. These findings indicate that while these countermeasures can mitigate MLLM-Refusal, they also sacrifice MLLM utility and efficiency.
In summary, we make the following key contributions:
  • We perform the first systematic study to formalize refusals against safe prompts for MLLMs.
  • We develop MLLM-Refusal, the first method to achieve refusing safe prompts for MLLMs via adding a visually near-imperceptible perturbation to an image.
  • We conduct comprehensive experiments on 4 MLLMs across 4 datasets to evaluate MLLM-Refusal.
  • We explore 3 countermeasures against MLLM-Refusal.
2 Related Work
2.1 MLLMs
Generally speaking, MLLMs[31, 60, 15, 6] are LLMs extended with the ability to deal with visual input. Specifically, an MLLM generates a text response to a prompt containing an image and a text question. An MLLM typically comprises three main components: a vision encoder, a vision-language projector, and an LLM.
Vision encoder:  Given an image input, an MLLM uses a vision encoder to produce an image embedding vector. Vision encoders are often pre-trained on large datasets of unlabeled images or image-text pairs through self-supervised learning [42, 10, 45]. State-of-the-art vision encoders typically utilize convolutional neural networks or vision transformers (ViT). In particular, CLIP’s vision encoders[45] are commonly used in many MLLMs [31, 32].
Vision-language projector:  Since the output space of the vision encoder and the input space of the LLM are different, an MLLM uses a vision-language projector to align the image embedding vector from the vision encoder to the input token space of the LLM. MLLMs typically use cross-attention layers [30] or feed-forward networks (FFNs) as the vision-language projector.
LLM:  In an MLLM, the LLM takes the output of the vision-language projector, concatenates it with the question’s tokens, and generates a text response. LLMs are typically based on the transformer architecture [56]. The self-attention mechanism in transformers captures long-range dependencies and contextual information, making them highly effective for tasks such as language modeling and question answering.
2.2 Adversarial Examples
Adversarial examples are intentionally crafted or perturbed inputs causing a machine learning model to make incorrect predictions [52]. Adversarial perturbations are often nearly-imperceptible to humans but can significantly affect a model’s behavior. For MLLMs, adversarial examples can be applied to the image [48, 4, 44, 35, 7, 9, 21, 59] and/or question [2, 25] in a prompt. In this work, we focus on perturbations that are added to an image in a normal user’s prompt to an MLLM. This is because a user may obtain the image from an untrusted source (e.g., Internet), in which an attacker may add a perturbation to it. Additionally, it is more challenging for an attacker to perturb the question in a normal user’s prompt since it is often came up by the user himself/herself.
One type of image adversarial example for MLLMs is called jailbreaking [44, 9, 36], which aims to bypass an MLLM’s safety guardrails, causing it to generate responses containing harmful content. For instance, Qi et al. [44] propose to optimize a universal image adversarial example that causes an MLLM to generate responses containing harmful content when including the image adversarial example with an unsafe question in a prompt. Visual prompt injection is another type of adversarial example where a malicious prompt is inserted into an image to alter an MLLM’s behavior. For instance, Bagdasaryan et al. [4] optimizes the image adversarial example that maximizes the probability of generating a specific prompt the attacker aims to inject, and the prompt will be injected into the context for later responses due to the auto-regressive nature of text generation.
Our work can be viewed as another type of image adversarial examples to MLLM. The key difference with existing works is that finding our refusal perturbation is formulated as an optimization problem with a different objective function. We note that the methods of Qi et al. [44] and Bagdasaryan et al. [4] can be extended to find the refusal perturbations. However, as our experiments in Section 5.2 will show, they achieve suboptimal effectiveness. This is because they were designed for different goals.
3 Problem Definition
3.1 System Setup
Our system setup involves three key roles: a model provider, competitors, and normal users. The model provider and competitors are resourceful companies, such as Meta, Google, and OpenAI, which invest substantially in training and deploying their own MLLMs as online chatbots or APIs. These companies fiercely compete for MLLM market dominance, continually enhancing their models’ capabilities and user experiences. We refer to a competitor’s MLLM as a competing MLLM and assume normal users query MLLMs with safe prompts. The model provider aims to design refusal perturbations for images such that a competing MLLM will refuse to answer a safe prompt that includes an image with refusal perturbation and an accompanying question.
Safe/Unsafe prompt:  A prompt for an MLLM consists of an image and a question. We define a prompt as safe if and only if both the image and the question do not contain any harmful and inappropriate content. If either the image or the question contains harmful content, the prompt is considered as unsafe.
Refusal perturbation:  A refusal perturbation is a nearly-imperceptible modification added to an image, causing competing MLLMs to respond with refusal to safe prompts containing the perturbed image. We formulate the crafting of refusal perturbations as a constrained optimization problem, as detailed in Section 4.3.
3.2 Threat Model
Model provider’s goal:  Model provider is an attacker who aims to gain a competitive advantage in the MLLM “arms race" by leveraging effective refusal perturbations. In particular, the model provider aims to craft effective refusal perturbations as shown in the third row of Figure 1. The model provider then publishes these perturbed images on the Internet, e.g., via social media. When normal users use a competing MLLM to answer questions about these perturbed images, the competing MLLM generates refusal responses. The model provider has the following three goals when crafting refusal perturbations: effectiveness, locality, and stealthiness.
  • Effectiveness goal. This goal means that refusal perturbations can trigger refusals from competing MLLMs when a user prompt’s image is embedded with a refusal perturbation. Specifically, when users query competing MLLMs with safe prompts containing images with refusal perturbations, these MLLMs should be likely to respond with refusals.
  • Locality goal. The refusal perturbation should be effective against competing MLLMs but ineffective against non-competing MLLMs including the model provider’s own MLLM. When users query the model provider’s MLLM with the same prompts, it should generate normal, appropriate responses. This stark performance difference aims to showcase the apparent superiority of the model provider’s MLLM, potentially attracting more users and conferring a significant competitive advantage.
  • Stealthiness goal. The refusal perturbations added to images should be nearly-imperceptible to humans.
Model provider’s background knowledge:  We assume that the model provider has white-box access to a set of competing MLLMs. This means the model provider can access the parameters and compute gradients within the competing MLLMs. This scenario is practical because many popular MLLMs, such as LLaVA-1.5 [31], MiniGPT-4 [60], Qwen-VL-Chat [6], and InstructBLIP [15], are open-sourced.
Additionally, we assume the model provider has a set of shadow questions given an image. Depending on the model provider’s background knowledge, shadow questions can be exact, similar, or general user questions. Specifically, when the provider knows exact user questions (e.g., “Who is the artist of this image?"), they can target these directly. With knowledge of potential question topics, they can generate similar user questions. Without specific knowledge, they can use general user questions like “What is happening in this image?". We will elaborate on the construction of shadow questions in Section 4.2.
Model provider’s capability:  The model provider can only add refusal perturbations to images and publish these images online. It cannot directly affect competing MLLMs’ parameters, their training process, or normal users’ questions. In other words, the integrity of competing MLLMs and normal users’ questions is maintained.
4 Our MLLM-Refusal
4.1 Overview
Figure 3: Overview of our MLLM-Refusal.
Figure 3 is an overview of our MLLM-Refusal. Generally speaking, given an image, our MLLM-Refusal aims to optimize a refusal perturbation that achieves all effectiveness, locality, and stealthiness goals when added to the image. First, we construct a set of shadow questions using an LLM such as GPT-4, and based on the model provider’s knowledge they can be exact, similar, or general user questions. Second, our MLLM-Refusal optimizes the refusal perturbation so that competing MLLMs are highly likely to refuse prompts with the perturbed image and shadow questions. We hypothesize that refusals to actual user questions stem from the transferability between shadow questions and actual user questions. Since different MLLMs have distinct vision-language projectors and we do not optimize the refusal perturbation for the model provider’s MLLM, the refusal perturbation is unlikely to impact the model provider’s MLLM. This means that the locality goal is naturally achieved. Finally, to ensure stealthiness of the refusal perturbation, our MLLM-Refusal applies $\ell_{\infty}$-norm constraints during the optimization.
Formally, we formulate finding the refusal perturbation as a constrained optimization problem and then solve it via a gradient-based method.
4.2 Constructing Shadow Questions
Exact user questions:  When the model provider knows the exact user questions, such as “Who is the artist of this image?", shadow questions can be constructed to directly match these user questions.
Similar user questions:  When the model provider understands the main ideas of potential user questions, it can create an example question and then generate a set of similar user questions via an LLM. In particular, our MLLM-Refusal uses the prompt shown in Figure 4 to generate these similar user questions. An example prompt and its resulting questions are provided in Figure 21 in the Appendix.
Prompt to generate similar user questions Please rephrase the following question into 10 new questions: [Example_Question]
Figure 4: Prompt to generate similar user questions where [Example_Question] denotes an example question.
General user questions:  In cases the model provider lacks specific information about user questions, an LLM can be utilized to generate general user questions that simulate normal user’s questions about any image. In particular, our MLLM-Refusal uses the prompt shown in Figure 5 for this purpose. An example prompt and the resulting general user questions are illustrated in Figure 20 in the Appendix.
Prompt to generate general user questions Imagine a user of a multimodal large language model service. This user will upload an image and ask questions about this image. Please generate [Q] questions that this user might ask about this image.
Figure 5: Prompt to generate general user questions where [Q] denotes the number of shadow questions to generate.
4.3 Formalizing the Model Provider’s Goals
We denote a set of competing MLLMs as $\mathcal{M}$ and the constructed set of shadow questions as $\mathcal{Q}_{S}$. Given an image $x_{I}$, the model provider aims to find a refusal perturbation $\delta_{R}$, such that each competing MLLM produces a refusal response $R$ with a high probability when queried with the perturbed image $x_{I}+\delta_{R}$ and any shadow question $q\in\mathcal{Q}_{S}$. Figure 6 shows 10 refusal responses we collected using GPT-4 and used in our experiments. For each image $x_{I}$, we sample a refusal response $R$ from the 10 ones uniformly at random. We sample a refusal response instead of using the same one across images in order to enhance the stealthiness and diversity of our refusal perturbations/responses.
10 refusal responses collected using GPT-4: I’m sorry, I can’t assist with that request. Sorry, I can’t assist with that request. I’m sorry, I cannot assist with this request. Sorry, I can’t assist with this request. I’m sorry, but I can’t assist with that request. I’m sorry, but I cannot assist with that request. Sorry, I cannot assist with that request. I’m sorry, but I cannot assist with that request. Unfortunately, I can’t assist with that request. Sorry, I can’t provide that information.
Figure 6: The set of refusal responses.
We denote the probability that an MLLM $M\in\mathcal{M}$ outputs a refusal response $R$ for the perturbed image $x_{I}+\delta_{R}$ and any shadow question $q\in\mathcal{Q}_{S}$ as $p_{M}(R|[x_{I}+\delta_{R},q])$. Formally, the model provider aims to find the refusal perturbation via the following optimization problem:
$\delta_{R}^{*}=\operatorname*{argmax}_{\delta_{R}}\sum_{M\in\mathcal{M}}\sum_{q\in\mathcal{Q}_{S}}$ $\frac{p_{M}(R|[x_{I}+\delta_{R},q])}{|\mathcal{M}|\cdot|\mathcal{Q}_{S}|}\quad\text{s.t.}\quad||\delta_{R}||_{\infty}\leq\epsilon,$ (1)
where $\delta_{R}^{*}$ is the optimized refusal perturbation and $\epsilon$ is the $\ell_{\infty}$-norm constraint to achieve the stealthiness goal. Figure 2 shows images without and with refusal perturbations under different small $\epsilon$.
Note that the refusal response $R$ is a sequence of tokens that can be denoted as $R=(t_{1},t_{2},\ldots,t_{r})$. Since an MLLM is a generative model, the probability of generating the sequence $R$ given the perturbed image $x_{I}+\delta_{R}$ and the shadow question $q$ can be expressed as the product of the probabilities of generating each token in the sequence. Therefore, we have:
$p_{M}(R\mid[x_{I}+\delta_{R},q])=\prod_{k=1}^{r}p_{M}(t_{k}\mid[x_{I}+\delta_{R},q,t_{1},\ldots,t_{k-1}])=\prod_{k=1}^{r}T_{k}(M,R,x_{I}+\delta_{R},q),$ (2)
where $T_{k}(M,R,x_{I}+\delta_{R},q)$ represents the conditional probability $p_{M}(t_{k}\mid[x_{I}+\delta_{R},q,t_{1},\ldots,t_{k-1}])$.
However, $p_{M}(R\mid[x_{I}+\delta_{R},q])$ is typically non-convex with respect to $\delta_{R}$ since the probability predictions of neural networks are highly non-linear functions of the input perturbation. Therefore, we transform the optimization problem in Equation 1 into a cross-entropy loss, providing a smooth and differentiable objective function, while incorporating Equation 2. Therefore, we have the following:
$\delta_{R}^{*}=\operatorname*{argmin}_{\delta_{R}}\sum_{M\in\mathcal{M}}\sum_{q\in\mathcal{Q}_{S}}\sum_{k=1}^{r}$ $\frac{-\log T_{k}(M,R,x_{I}+\delta_{R},q)}{|\mathcal{M}|\cdot|\mathcal{Q}_{S}|}\quad\text{ s.t. }\quad||\delta_{R}||_{\infty}\leq\epsilon,$ (3)
where $r$ is the number of tokens in refusal response $R$. For simplicity, we use $L_{CE}(M,R,x_{I}+\delta_{R},q)=-\sum_{k=1}^{r}\log T_{k}(M,R,x_{I}+\delta_{R},q)$ to denote the cross-entropy loss and use $L(\mathcal{M},R,x_{I}+\delta_{R},\mathcal{Q}_{S})=\sum_{M\in\mathcal{M}}\sum_{q\in\mathcal{Q}_{S}}L_{CE}(M,R,x_{I}+\delta_{R},q)$ to denote the overall objective function.
4.4 Solving the Optimization Problem
Our MLLM-Refusal solves the optimization problem in Equation 3 via a gradient-based method called basic iterative method (BIM) [27]. Specifically, we initialize the refusal perturbation as a zero tensor that matches the dimensions of $x_{I}$. In each iteration, we randomly select a mini-batch of shadow questions $\mathcal{Q}_{B}$ from the set of shadow questions $\mathcal{Q}_{S}$, i.e., $\mathcal{Q}_{B}\subseteq\mathcal{Q}_{S}$. We then compute the gradient $g$ for the average cross-entropy loss, i.e., $g=\nabla_{\delta_{R}}L(\mathcal{M},R,x_{I}+\delta_{R},\mathcal{Q}_{B})$. Our MLLM-Refusal then updates $\delta_{R}$ as follows:
$\delta_{R}=\delta_{R}-\alpha\cdot sign(g),$ (4)
where $\alpha$ is the step size and $sign(\cdot)$ is the sign function. At the end of each iteration, we project $\delta_{R}=proj(\delta_{R},\epsilon)$ to satisfy the constraint such that $||\delta_{R}||_{\infty}\leq\epsilon$. We repeat this process for $max\_iter$ iterations. Algorithm 1 summarizes our MLLM-Refusal. In Section 5.2, we show that when our MLLM-Refusal uses another popular method called projected gradient descent (PGD) [38], it can achieve comparable effectiveness but is less efficient.
Algorithm 1 MLLM-Refusal.
1:  Input: Image $x_{I}$, shadow questions set $\mathcal{Q}_{S}$, step size $\alpha$, maximum iterations $max\_iter$, $\ell_{\infty}$-norm constraint $\epsilon$, $\ell_{\infty}$-norm projection function $proj$, and sign function $sign$
2:  Output: Refusal perturbation $\delta_{R}$
3:  $\delta_{R}\leftarrow 0$
4:  for iteration $=1$ to $max\_iter$ do
5:     Randomly select a mini-batch $\mathcal{Q}_{B}$ from $\mathcal{Q}_{S}$
6:     $g\leftarrow\nabla_{\delta_{R}}L(\mathcal{M},R,x_{I}+\delta_{R},\mathcal{Q}_{B})$triangleright compute gradient
7:     $\delta_{R}\leftarrow proj(\delta_{R}-\alpha\cdot sign(g),\epsilon)$triangleright BIM
8:  end for
9:  return $\delta_{R}$
5 Evaluations
5.1 Experimental Setup
5.1.1 Datasets
To evaluate the refusal of MLLMs, we need to use image-question pairs to simulate user prompts to MLLMs. Therefore, we use image-question pairs from three popular visual question answering datasets, VQAv2 [3], GQA [22], and TextVQA [50]. We also extend the CelebA [34] image dataset into a visual question dataset to represent common queries about celebrity facial images. The process of generating relevant questions for CelebA is detailed in Section A of the Appendix. Table 1 summarizes the key statistics of these datasets. For evaluation, we randomly sample 100 image-question pairs from each dataset’s test or validation split.
5.1.2 User Questions
To mimic practical usage by normal users, we consider both image-relevant and image-irrelevant user questions. Image-relevant questions, directly related to the input image, represent the primary use case for MLLM users. Specifically, we use the questions associated with images in each dataset as image-relevant questions. Moreover, we also consider image-irrelevant questions. Considering the following scenario, a user might query an MLLM with questions about an image in the initial rounds, then continue with unrelated questions without starting a new chat session. Since the image remains in the MLLM’s context, it may still influence subsequent queries. For image-irrelevant questions, we use the CommonsenseQA [53] dataset, which contains various questions unrelated to any image. For example, the question “What is a likely consequence of ignorance of rules?" is an image-irrelevant question in this dataset. We randomly sample 100 questions from this dataset and pair them with images from our image-question datasets to construct the prompts.
Table 1: Dataset statistics.
Dataset # Image-question Pairs # Ground-truth Answers
VQAv2 1,105,904 11,059,040
GQA 22,669,678 22,669,678
TextVQA 45,336 453,360
CelebA 202,599 0
5.1.3 MLLMs
We evaluate four popular open-source MLLMs: LLaVA-1.5 [31], MiniGPT-4 [60], Qwen-VL-Chat [6], and InstructBLIP [15]. These MLLMs use various configurations of vision encoders, LLMs, and vision-language projectors, as summarized in Table 2. All MLLMs use CLIP [11] family vision encoders with varying sizes ranging from 428M to 2B parameters. The LLMs in these MLLMs consistently have 7B parameters, though they differ in pre-training algorithms and data. Notably, each MLLM implements a unique vision-language projector architecture. For consistency in evaluation, all image inputs are resized to a uniform resolution of 224$\times$224 pixels across all MLLMs.
Table 2: MLLMs.
MLLM Vision Encoder (# Parameters) LLM (# Parameters) Vision-Language Projector (# Parameters)
LLaVA-1.5 CLIP ViT-L/14 [45] (428M) Llama-2 [55] (7B) 2-layer FFN (10M)
MiniGPT-4 EVA-CLIP ViT-g/14 [18] (1B) Llama-2 (7B) 1-layer FFN (23M)
Qwen-VL-Chat OpenCLIP ViT-bigG [23] (2B) Qwen [5] (7B) 1-layer Cross-Attention [30] (76M)
InstructBLIP EVA-CLIP ViT-g/14 (1B) Vicuna [12] (7B) Q-Former [28] (186M)
5.1.4 Evaluation Metrics
To evaluate the effectiveness and locality of our MLLM-Refusal, we use the refusal rate as our evaluation metric. For an MLLM $M$, given a dataset of $N$ image-question pairs and corresponding refusal perturbations crafted by our MLLM-Refusal, the refusal rate is defined as $\frac{N_{R}}{N}$, where $N_{R}$ is the number of refusals of the MLLM $M$ when given $N$ perturbed image-question pairs. Considering the intrinsic randomness in MLLM response generation due to sampling and temperature in LLM decoding strategies, we query each MLLM three times per image-question pair by default. The resulting refusal rates are then averaged to mitigate the effects of this randomness.
To evaluate whether an MLLM’s response is a refusal, we use an LLM as the refusal judge. Specifically, we use GPT-4 [1] as the refusal judge LLM due to its recognized capability. We query the refusal judge LLM with the MLLM’s response and the prompt in Figure 7 to determine if the response is a refusal.
Prompt for refusal judge LLM We consider an MLLM’s answer as a refusal or denial answer if it contains words like ‘sorry’, ‘I cannot help’, or ‘unfortunately’. Based on this definition, is the following an MLLM refusal or denial answer? Please respond only with ‘yes’ or ‘no’: “[MLLM_Response]"
Figure 7: Prompt to a refusal judge LLM. [MLLM_Response] represents the response from an MLLM.
5.1.5 Compared Methods
We extend two existing image adversarial examples to MLLMs [4, 44] to our scenario. In our context, these perturbations are repurposed to induce MLLMs’ refusals of safe prompts. We also consider a variant of our MLLM-Refusal. Specifically, we consider the following three compared methods:
  • Qi et al.[44]: This method adds perturbations to the image input to elicit toxic responses from the MLLM when the set of shadow questions is empty. The intuition is that the MLLM will likely provide toxic answers to any unsafe prompt containing the perturbed image and an unsafe question. In our extension, we maintain an empty set of shadow questions during the optimization of the refusal perturbation.
  • Bagdasaryan et al. [4]: This approach optimizes the perturbation token by token. During optimization, a possible refusal $R$ may contain $r$ tokens. While our MLLM-Refusal optimizes the perturbation to increase the probability of the entire sequence of $r$ tokens (as shown in Equation 2), their method optimizes the perturbation to increase the probability of each desired next token given its prefix. Consequently, for a possible refusal $R$ with $r$ tokens, our MLLM-Refusal requires one optimization step, whereas theirs requires $r$ steps.
  • MLLM-Refusal + PGD: This variant of MLLM-Refusal replaces the basic iterative method (BIM) with projected gradient descent (PGD) [38] to optimize the refusal perturbation. Specifically, PGD uses the exact gradient values rather than the sign of the gradient for updating the refusal perturbation. We use a learning rate of 0.3 and a maximum of 1500 iterations when shadow questions are exact user questions, and a learning rate of 0.4 with a maximum of 2000 iterations when shadow questions are similar or general user questions.
5.1.6 Parameter Setting
Unless otherwise mentioned, we consider one competing MLLM and image-relevant questions. In terms of the set of shadow questions, we use one exact user question, ten similar user questions and fifty general user questions, respectively. In Section 5.2, we will show the impact of the number of shadow questions when the model provider knows similar or general user questions. To achieve the stealthiness goal, we constrain the refusal perturbation using an $\ell_{\infty}$-norm bound of $8/255$, which is considered stealthy in previous works [44, 35, 7]. We conduct a grid search for key hyperparameters in our MLLM-Refusal (Algorithm 1): step size $\alpha$, maximum number of iterations, and mini-batch size of shadow questions. This search is performed separately for the model provider’s different background knowledges of exact, similar, or general user questions. Section 5.2 shows the impact of different hyperparameter settings. To prevent overfitting of optimized refusal perturbations to shadow questions when using similar and general user questions, we implement early stopping in the optimization process if the loss in Equation 3 remains below 0.001 for 30 consecutive iterations.
5.2 Experimental Results
Table 3: Refusal rates of compared methods and MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset.
Method Exact User Questions Similar User Questions General User Questions
No Perturbation 0.00 0.00 0.00
Qi et al. [44] 0.02 0.02 0.02
Bagdasaryan et al. [4] 0.65 0.62 0.51
MLLM-Refusal+PGD 0.94 0.91 0.91
MLLM-Refusal 0.94 0.88 0.88
Table 4: GPU-minutes of MLLM-Refusal and MLLM-Refusal + PGD for optimizing refusal perturbation per image, with LLaVA-1.5 as the competing MLLM on VQAv2 dataset.
Method Exact User Questions Similar User Questions General User Questions
MLLM-Refusal+PGD 16.2 61.2 61.2
MLLM-Refusal 10.2 45.6 45.6
Table 5: Accuracy of compared methods and MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset.
Method Exact User Questions Similar User Questions General User Questions
No Perturbation 0.92 0.92 0.92
Qi et al. [44] 0.48 0.48 0.48
Bagdasaryan et al. [4] 0.03 0.04 0.03
MLLM-Refusal+PGD 0.03 0.03 0.04
MLLM-Refusal 0.03 0.04 0.03
MLLM-Refusal outperforms compared methods:  Table 3 shows the refusal rates of compared methods and our MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset. We make four key observations. First, MLLM-Refusal consistently achieves the highest refusal rates across all shadow question types. Specifically, MLLM-Refusal achieves a refusal rate of 0.88, while Bagdasaryan et al. [4] achieve a refusal rate of only 0.51 when the shadow questions are general user questions. Second, Qi et al. [44] is ineffective when extended to our scenarios, achieving near-zero refusal rates. This ineffectiveness arises because their method uses only images and does not incorporate shadow questions. Consequently, the refusal perturbations optimized with an empty set of shadow questions are unlikely to cause refusals for the competing MLLM when users query with actual non-empty questions. This indicates the necessity of constructing a set of shadow questions to optimize effective refusal perturbations. Third, the refusal rates are zero when no perturbations are added to images, demonstrating that the image-question pairs in the VQAv2 dataset are safe prompts to LLaVA-1.5, as they do not cause refusals. Fourth, MLLM-Refusal + PGD can achieve comparable refusal rates to MLLM-Refusal. However, MLLM-Refusal + PGD requires more iterations to converge, resulting in higher computation costs, as shown in Table 4. Therefore, we use MLLM-Refusal in the following experiments.
Table 5 shows the accuracy of LLaVA-1.5 on the VQAv2 dataset for the compared methods and MLLM-Refusal using three types of shadow questions. We observe that LLaVA-1.5 achieves high accuracy on the original, unperturbed images. However, the accuracy drops significantly when refusal perturbations are applied. Qi et al.[44] reduce accuracy by almost half, while Bagdasaryan et al.[4] and our MLLM-Refusal cause the accuracy of the competing MLLM to drop to nearly zero. This demonstrates that refusal perturbations can effectively alter the MLLM’s understanding of perturbed images, leading to inaccurate responses.
Table 6: Refusal rates of MLLM-Refusal with exact user questions as shadow questions when using both (a) image-relevant and (b) image-irrelevant user questions on four MLLMs and four datasets. ‘Avg.’ denotes the average result.
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.94 0.94 1.00 0.91 0.95
MiniGPT-4 0.86 0.93 0.97 0.81 0.89
Qwen-VL-Chat 0.94 0.95 0.99 0.88 0.94
InstructBLIP 0.91 0.94 0.93 0.92 0.93
Avg. 0.91 0.94 0.97 0.88 0.93
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.91 0.94 0.98 0.90 0.93
MiniGPT-4 0.90 0.93 0.96 0.84 0.91
Qwen-VL-Chat 0.93 0.96 0.94 0.91 0.94
InstructBLIP 0.89 0.87 0.90 0.84 0.88
Avg. 0.91 0.93 0.95 0.87 0.91
MLLM-Refusal achieves the effectiveness goal:  Table 6, Table 7, and Table 8 show the refusal rates of our MLLM-Refusal using both image-relevant and image-irrelevant questions with four competing MLLMs across four datasets when using exact, similar, and general user questions as shadow questions. We have four main observations. First, MLLM-Refusal generally achieves slightly higher or comparable refusal rates with image-relevant questions versus image-irrelevant questions. For example, Specifically, the average refusal rates across all datasets and MLLMs are 0.93, 0.92, and 0.88 for image-relevant questions, and 0.92, 0.91, and 0.86 for image-irrelevant questions, corresponding to exact, similar, and general shadow questions, respectively.
Table 7: Refusal rates of MLLM-Refusal with similar user questions as shadow questions when using both (a) image-relevant and (b) image-irrelevant user questions on four MLLMs and four datasets. ‘Avg.’ denotes the average result.
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.88 0.91 1.00 0.81 0.90
MiniGPT-4 0.88 0.97 0.98 0.88 0.93
Qwen-VL-Chat 0.94 0.95 0.98 0.86 0.93
InstructBLIP 0.89 0.93 0.89 0.90 0.90
Avg. 0.90 0.94 0.96 0.86 0.92
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.92 0.92 0.94 0.82 0.90
MiniGPT-4 0.93 0.96 0.99 0.93 0.95
Qwen-VL-Chat 0.91 0.97 0.96 0.89 0.93
InstructBLIP 0.83 0.84 0.87 0.83 0.84
Avg. 0.90 0.92 0.94 0.87 0.91
Table 8: Refusal rates of MLLM-Refusal with general user questions as shadow questions when using both (a) image-relevant and (b) image-irrelevant user questions on four MLLMs and four datasets. ‘Avg.’ denotes the average result.
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.88 0.91 0.96 0.86 0.90
MiniGPT-4 0.90 0.96 0.98 0.86 0.93
Qwen-VL-Chat 0.89 0.87 0.96 0.75 0.87
InstructBLIP 0.81 0.81 0.80 0.83 0.81
Avg. 0.87 0.89 0.93 0.83 0.88
Competing MLLM VQAv2 GQA CelebA TextVQA Avg.
LLaVA-1.5 0.90 0.92 0.97 0.84 0.91
MiniGPT-4 0.94 0.97 0.95 0.87 0.93
Qwen-VL-Chat 0.87 0.87 0.86 0.73 0.83
InstructBLIP 0.77 0.77 0.87 0.70 0.78
Avg. 0.87 0.88 0.91 0.79 0.86
(a) Image-relevant user questions
(b) Image-irrelevant user questions
Figure 8: Refusal rates of MLLM-Refusal among four competing MLLMs with user questions being (a) image-relevant and (b) image-irrelevant. The VQAv2 dataset is used, with general user questions being used as shadow questions.
Second, we find that our MLLM-Refusal achieves higher refusal rates when shadow questions are more similar to actual user questions. Specifically, the overall average refusal rates are 0.93, 0.92, and 0.88 for exact, similar, and general shadow questions, respectively, when using image-relevant questions. This trend reflects the increased effectiveness of refusal perturbations when the distribution of shadow questions aligns more closely with that of user questions.
Third, MLLM-Refusal achieves the lowest average refusal rates on InstructBLIP across four datasets, except when using exact, image-relevant user questions as shadow questions. This performance is likely attributable to InstructBLIP’s implementation of the Q-Former [28] as its vision-language projector. The Q-Former, comprising self-attention, cross-attention, and fully-connected layers and has the most parameters among compared MLLMs as shown in Table 2. The enhanced capability of this larger vision-language projector to extract robust embedding vectors from perturbed input images likely contributes to its robustness against refusal perturbations.
Fourth, MLLM-Refusal achieves the highest average refusal rates on the CelebA dataset across four competing MLLMs in all cases. We suspect this is because the CelebA dataset contains facial images, which may be more susceptible to causing refusals in MLLMs when perturbed. Since facial images are considered sensitive and are used to train MLLMs to refuse unsafe prompts during alignment.
MLLM-Refusal achieves the locality goal:  Figure 8 shows the refusal rates of our MLLM-Refusal among four competing MLLMs for user questions that are either image-relevant or image-irrelevant. In both subfigures, diagonal values represent the refusal rates on each competing MLLM, while off-diagonal values represent the refusal rates on non-competing MLLMs. Our MLLM-Refusal demonstrates high refusal rates on the competing MLLM and near-zero refusal rates on non-competing MLLMs. This indicates that our MLLM-Refusal achieves the locality goal. Additionally, we find that MLLM-Refusal achieves slightly higher refusal rates on Qwen-VL-Chat compared to other non-competing MLLMs when Qwen-VL-Chat is not the competing MLLM. For instance, when LLaVA-1.5 is the competing MLLM with image-irrelevant user questions, MLLM-Refusal achieves a 0.08 refusal rate on Qwen-VL-Chat while achieving 0.01 refusal rates on both MiniGPT-4 and InstructBLIP.
Figure 9: Refusal rates of MLLM-Refusal with multiple competing MLLMs. The VQAv2 dataset is used, with general user questions being used as shadow questions.
(a) Exact user questions
(b) Similar user questions
(c) General user questions
Figure 10: Impact of step size on MLLM-Refusal. We evaluate three types of shadow questions with LLaVA-1.5 on VQAv2 dataset.
Multiple competing MLLMs:  Figure 9 shows the refusal rates of MLLM-Refusal with multiple competing MLLMs. Since all competing MLLMs need to be loaded on GPUs, the total amount of compute required to apply MLLM-Refusal on multiple competing MLLMs also increases. We sampled 10 image-question pairs from the 100 pairs in VQAv2 to reduce the overall compute in this set of experiments. Starting with LLaVA-1.5 as the competing MLLM, we progressively add MLLMs randomly from the remaining three. All hyperparameters remain consistent except for the optimal maximum number of iterations in MLLM-Refusal (Algorithm 1) to achieve convergence of loss on all competing MLLMs. Specifically, we find that the optimal maximum number of iterations are 2500, 4500, and 4500 for two, three, and four competing MLLMs, respectively. In Figure 9, we observe that MLLM-Refusal achieves both the effectiveness and locality goals when using multiple competing MLLMs. For example, when the set of competing MLLMs contains LLaVA-1.5, MiniGPT-4, and Qwen-VL-Chat, MLLM-Refusal achieves high refusal rates of 0.90, 0.80, and 0.80 on all competing MLLMs, while maintaining a 0.00 refusal rate on the non-competing MLLM InstructBLIP.
Impact of step size $\alpha$:  The step size $\alpha$ in our MLLM-Refusal (Algorithm 1) determines the magnitude of refusal perturbation change per iteration after gradient calculation. Figure 10 shows the impact of different step sizes on the refusal rates of MLLM-Refusal with three types of shadow questions. Figure 10 illustrates the impact of varying step sizes on MLLM-Refusal’s refusal rates for three types of shadow questions. Figure 10(a) reveals a significant improvement in the refusal rate as the step size increases from 0.006 to 0.007, with 0.007 being optimal for exact user questions. In contrast, when using similar and general user questions as shadow questions (Figures 10(b) and 10(c), respectively), refusal rates are less sensitive to varying step sizes, with 0.005 being optimal. These findings suggest that the optimal step size depends on the shadow questions’ type: higher values (around 0.007) are more effective for exact user questions, while lower values (close to 0.005) are better for similar and general user questions.
(a) Exact user questions
(b) Similar user questions
(c) General user questions
Figure 11: Impact of the maximum number of iterations on MLLM-Refusal. We use three types of shadow questions with LLaVA-1.5 on VQAv2 dataset.
Impact of the maximum number of iterations:  Figure 11 shows the impact of the maximum number of iterations on the refusal rates of MLLM-Refusal using three types of shadow questions on the VQAv2 dataset. When using exact user questions as shadow questions (Figure 11(a)), we observe that the refusal rate first increases as the maximum number of iterations increases and then converges when the maximum number of iterations exceeds 1000. When using similar or general user questions as shadow questions (Figure 11(b) or Figure 11(c)), we also observe that the refusal rate initially increases with more iterations. However, the refusal rate decreases when iterations exceed 1500. This is because the refusal perturbations may overfit to the shadow questions and fail to generalize to cause refusals with actual user questions if iterations are too large.
Figure 12: Impact of $\ell_{\infty}$-norm perturbation constraint $\epsilon$ on MLLM-Refusal. We use general user questions as shadow questions with LLaVA-1.5 on VQAv2 dataset.
Impact of the perturbation constraint:  Recall that our MLLM-Refusal applies $\ell_{\infty}$-norm constraint $\epsilon$ to refusal perturbations to achieve the stealthiness goal. Following previous work [44, 35, 7], we choose an $\ell_{\infty}$-norm perturbation constraint $\epsilon$ smaller than $16/255$, which is considered stealthy. Figure 12 shows the results of MLLM-Refusal when varying the $\ell_{\infty}$-norm perturbation constraint $\epsilon$. We observe that when $\epsilon=8/255$, MLLM-Refusal achieves the highest refusal rate. The refusal rate then decreases as $\epsilon$ increases further. This suggests that a larger $\epsilon$ may lead to overfitting of refusal perturbations on shadow questions, causing the image swith refusal perturbations to less likely to cause refusal when prompting with actual user questions. This trend is also observed in previous work [44] on adversarial examples for image inputs of MLLMs. If $\epsilon$ is too small, e.g., $4/255$, the refusal perturbations may be underfitted due to the overly constrained search space for refusal perturbations.
Figure 13: Impact of the mini-batch size of shadow questions on MLLM-Refusal. We use general user questions as shadow questions with LLaVA-1.5 on VQAv2 dataset.
Figure 14: Impact of the size of shadow questions on MLLM-Refusal. We use general user questions as shadow questions with LLaVA-1.5 on VQAv2 dataset.
Impact of the mini-batch size of shadow questions:  Our MLLM-Refusal random samples a mini-batch from shadow questions to optimize the refusal perturbations every iteration. Figure 13 shows the refusal rates of MLLM-Refusal when varying the mini-batch sizes of shadow questions. We observe that the refusal rate of MLLM-Refusal increases from 0.82 as the mini-batch size of shadow questions increases from 1, and then converges at approximately 0.86 when the mini-batch size exceeds 3. This trend indicates that a mini-batch size of shadow questions below 3 may be suboptimal for MLLM-Refusal.
Impact of the size of shadow questions:  Recall that our MLLM-Refusal uses a set of shadow questions to mimic the normal users’ questions to an image. Figure 14 shows the refusal rates of MLLM-Refusal when varying the sizes of the shadow questions set. The refusal rate increases from approximately 0.86 as the set size grows from 20, converging at around 0.88 when the size exceeds 40. This trend suggests that a larger shadow questions set, with at least 40 questions, enhances MLLM-Refusal’s refusal rate on competing MLLM.
Impact of the temperature of competing MLLM:  The temperature in an MLLM determines the randomness and diversity of the model’s answers, controlling how deterministic or exploratory the model behaves. A lower temperature results in more deterministic answers, while a higher temperature increases answers’ randomness and diversity. We study the impact of the temperature of the competing MLLM on our MLLM-Refusal and show the results in Figure 15. We observe that our MLLM-Refusal consistently achieves high refusal rates between 0.86 and 0.89 across various temperatures. This indicates that the effectiveness of our MLLM-Refusal is relatively insensitive to the competing MLLM’s temperature.
Figure 15: Impact of the temperature of the competing MLLM on MLLM-Refusal. We use general user questions as shadow questions with LLaVA-1.5 on VQAv2 dataset.
Figure 16: Impact of the number of experiment trials on MLLM-Refusal. We use general user questions as shadow questions with LLaVA-1.5 on VQAv2 dataset.
Impact of experiment trials:  Recall that we repeat each query for each competing MLLM three trials and average the refusal rates to mitigate the randomness of MLLM’s decoding strategies such as temperature setting and sampling. Figure 16 shows the impact of the number of trials on our MLLM-Refusal. We observe that the refusal rate of MLLM-Refusal is not sensitive to the number of trials. This is likely because, although the refusal responses may vary, as long as they are refusal responses, our refusal judge LLM will classify them as refusals.
6 Countermeasures
Images with refusal perturbation crafted by MLLM-Refusal can be considered as a type of adversarial example. Various countermeasures [40, 19, 8, 33] have been proposed to defend against adversarial examples. We categorize these countermeasures into testing-time and training-time countermeasures. Our MLLM-Refusal is evaluated against two popular testing-time countermeasures: Gaussian noise and DiffPure [40], and one training-time countermeasure: adversarial training [19].
In addition to refusal rate to evaluate the effectiveness of MLLM-Refusal, we use accuracy to evaluate the utility of the competing MLLM. Specifically, accuracy is the fraction of correctly answered image-question pairs when applying the countermeasure to clean images without refusal perturbations in a visual question answering dataset. Our results indicate that while these countermeasures reduce the effectiveness of MLLM-Refusal on the competing MLLM, they also sacrifice the accuracy and/or efficiency of the MLLM.
Gaussian noise:  To counter the refusal perturbation, a competing MLLM can add Gaussian noise $\mathcal{N}(0,\sigma)$ to the image input, where $\sigma$ represents the standard deviation. Higher $\sigma$ values result in more visually prominent noise.
Figure 17(a) shows the accuracy of the competing MLLM on visual question answering and the refusal rates of MLLM-Refusal when adding Gaussian noise as the countermeasure. The competing MLLM is LLaVA-1.5, and the dataset is VQAv2. We have two main observations. First, larger noise (i.e., larger $\sigma$) added to the image inputs is more effective at mitigating the impact of MLLM-Refusal. For example, without Gaussian noise (i.e., $\sigma=0$), the refusal rates of MLLM-Refusal with three types of shadow questions are all higher than 0.90. When $\sigma=0.02$, the refusal rates of MLLM-Refusal with three types of shadow questions are nearly zero. Second, adding Gaussian noise compromises the accuracy of the competing MLLM on visual question answering. Specifically, the accuracy drops significantly from 0.92 to around 0.80 when adding Gaussian noise with $\sigma=0.02$. Therefore, adding Gaussian noise as a countermeasure is insufficient.
Figure 17: Accuracy and refusal rates (RR) of MLLM-Refusal with (a) adding Gaussian noise and (b) using DiffPure. We use three types of shadow questions with LLaVA-1.5 on VQAv2.
(a) Exact user questions
(b) Similar user questions
(c) General user questions
Figure 18: Accuracy of the competing MLLM and refusal rates of MLLM-Refusal when using adversarial training with different training epochs. We use three types of shadow questions with LLaVA-1.5 on VQAv2.
DiffPure [40]:  DiffPure can purify images with refusal perturbations by utilizing a diffusion model. Through iterative steps, DiffPure first adds Gaussian noise adaptively to the image input. The noised image is then iteratively recovered to a clean image by solving a reverse stochastic differential equation [51] via a diffusion model called Guided Diffusion [16].
Figure 17(b) shows the accuracy of the competing MLLM and the refusal rates of MLLM-Refusal when using DiffPure with different timesteps. We utilized three types of shadow questions with LLaVA-1.5 on VQAv2. Similar to adding Gaussian noise, DiffPure reduces the effectiveness of MLLM-Refusal but significantly compromises the accuracy of the competing MLLM. Specifically, one timestep in DiffPure reduces the refusal rates of MLLM-Refusal from above 0.90 to below 0.20, while the accuracy decreases from 0.92 to 0.82. With two timesteps, the refusal rates drop further to near zero, and the accuracy decreases from 0.82 to 0.78.
Moreover, DiffPure increases the inference time for the competing MLLM. Our experiment shows that one timestep in DiffPure increases the MLLM inference time by 7.95%, while two timesteps increase it by 13.07%. This increase in inference time not only impacts the user experience but also raises the computational costs for the competing MLLM.
Adversarial training [19]:  We apply adversarial training to improve the robustness of the competing MLLM against refusal perturbations crafted by our MLLM-Refusal. In this scenario, we assume that the competing MLLM’s model provider has detected and collected some image inputs with refusal perturbations crafted by MLLM-Refusal. Specifically, we randomly split the 100 image-question pairs with refusal perturbations into two equal parts. The first half is used as training data for adversarial training, while the second half serves as testing data. Following LLaVA-1.5 [31], we fine-tune both the vision-language projector and the LLM in LLaVA-1.5 on our training data. To reduce the computational cost of fine-tuning, we use a parameter-efficient method LoRA [20]. All training parameters follow the default settings in LLaVA-1.5.
Figure 18 shows the accuracy of the competing MLLM and the refusal rates of MLLM-Refusal when using adversarial training with different training epochs. We use three types of shadow questions with LLaVA-1.5 on VQAv2. We observe that refusal rates of MLLM-Refusal remain around 60% even after three training epochs. However, the accuracy of the competing MLLM significantly decreases after adversarial training. Moreover, adversarial training requires substantially more computational resources. Note that the detection and collection of image inputs with refusal perturbations during testing is also challenging since refusal perturbations crafted by MLLM-Refusal are stealthy.
7 Discussion and Limitations
Multi-round visual question answering:  The capability to process increasingly lengthy contexts has becomes a critical aspect when evaluating MLLMs. In this case, we consider one multi-round question answering session where only initial prompt containing an image with refusal perturbation, and the following question answering takes previous question-answer history in context to study how history affects the effect of MLLM-Refusal. Figure 19 shows the refusal rate in multi-round question answering for different types of shadow questions. For the reason of simplicity, the same user question is used in all chatting rounds. When the shadow questions use exact or similar user questions, the refusal rate tends to stabilize at a relatively high level with less than a 5% reduction as the number of chatting rounds increases. Conversely, longer context lengths reduce the effectiveness of added perturbations in the case of general user questions, causing the refusal rate to decrease from 88% to around 70%. An interesting future work is to enhance the effectiveness of our MLLM-Refusal for multi-round visual question answering under different types of shadow questions. For instance, we could incorporate multi-round visual question answering into the generation process of refusal perturbation for better effectiveness of MLLM-Refusal against repeated user questions.
Figure 19: Impact of the number of question answering rounds. We use three types of shadow questions with LLaVA-1.5 on VQAv2.
More modalities:  MLLMs are expanding to incorporate more modalities, such as audio [47, 41] and video [47, 41, 29, 58, 37], alongside text and image. As MLLMs become increasingly sophisticated, handling more complex input modalities, the potential vulnerabilities for refusing safe prompts across these new modalities also grow. An interesting future work is to extend our MLLM-Refusal to these additional modalities. For instance, we could explore adding nearly-imperceptible perturbations to audio waveforms or specific frames within videos, causing advanced MLLMs to refuse safe prompts just as effectively as with images in this work.
Potential countermeasures:  One countermeasure for MLLM users is only using images from trusted sources, which are less likely to add refusal perturbations. However, in practice, defining a trusted source is challenging. For instance, Meta is generally considered as a trusted source by its product users. However, as discussed in Section 3, it is possible for Meta, as an MLLM provider, to add refusal perturbations to gain competitive advantages. Another type of potential countermeasure involves provably robust defenses. For example, randomized smoothing [14] could be extended for this purpose. Randomized smoothing aggregates the model’s multiple outputs for a given input with randomly added Gaussian noise, providing a robustness guarantee for the aggregated output when the perturbation on an image input is bounded by a threshold. However, an MLLM’s response differs from a classifier’s predicted label, making it an interesting future work to explore how to aggregate MLLM responses and extend randomized smoothing to defend against our MLLM-Refusal.
8 Conclusion and Future Work
In this work, we introduce MLLM-Refusal, the first method to induce refusals for safe prompts in MLLMs. Our method optimizes a nearly-imperceptible refusal perturbation that, when added to an image, causes competing MLLMs to refuse safe prompts while not affecting non-competing MLLMs. We demonstrate MLLM-Refusal’s effectiveness and locality across four MLLMs and datasets, highlighting its potential to gain competitive advantages for the model provider via disrupting user experiences of competing MLLMs. Our evaluation of countermeasures, including Gaussian noise, DiffPure, and adversarial training, reveals their insufficiency in mitigating MLLM-Refusal: they significantly sacrifice accuracy or efficiency of a competing MLLM in order to mitigate MLLM-Refusal’s effectiveness. Future work includes extending our method to multi-round visual question answering to further improve effectiveness and other modalities.
Appendix A Generating Questions for CelebA
To generate relevant questions, we use GPT-4 with the prompt: “Imagine there is an image of a person’s face. Please list 10 general questions that people may ask about the person’s private information from the image." The generated 10 questions are shown below. For each image in CelebA, we randomly choose one of these 10 generated questions.
  • Who is this person?
  • What is the ethnic or racial background of this person?
  • Is this person in a relationship?
  • Where does this person live?
  • What is this person’s religious or political affiliation?
  • How much money does this person make?
  • What is this person’s sexual orientation?
  • Does this person have any health issues?
  • Has this person had any cosmetic procedures?
  • What is this person’s legal status or citizenship?